OpenStudy (anonymous):

A collection of 30 coins worth $5.50 consists of nickels, dimes, and quarters. There are twice as many dimes as nickels. How many quarters are there?

OpenStudy (anonymous):

1.A collection of 30 coins worth $5.50 consists of nickels, dimes, and quarters. There are twice as many dimes as nickels. How many quarters are there? : The number of each coin equation: n + d + q = 30 : The $ amt equation: .05n + .10d + .25q = 5.50 : "There are twice as many dimes as nickels" equation: d = 2n : Substitute 2n for d in both equations n + 2n + q = 30 and .05n + .10(2n) + .25q = 5.50 Which is: 3n + q = 30 and .25n + .25q = 5.50 : Multiply the above equation by 4 and subtract it from 3n + q = 30 3n + 1q = 30 1n + 1q = 22 -------------- 2n + 0q = 8 n = 8/2 n = 4 nickels : Then we know that d = 2(4) = 8 dimes : Use the coins equation to find q: 4 + 8 + q = 30 q = 30 - 12 q = 18 quarters : Check solutions using the $ equation: .05(4) + .10(8) + .25(18) = .20 + .80 + 4.50 = 5.50